Coulomb's Law and Lattice Energy - Calculate amount of energy released

Calculate the amount of energy released in the formation of one mole of SrO bonds (not lattice energy). The radius of the strontium ion is 1.13 Å, and the radius ofthe oxide ion is 1.4 Å. Note that 1Å=10 to power -10 m
Need answer to be in angstroms.
Answer #1

2200kJ/mol
Answer #2
F=q*q/r*r (Coulomb's law)

=2*2/(1.6+1.26)*(10`-10)`2

=1.251*10`-35 kJ
answered by: weng
Answer #3

2.20X10^6 KJ/mol